# ALPINE: Unveiling the Planning Capability of Autoregressive Learning in Language Models

### v2

@@ Abstract
In this paper, we present the findings of our Project ALPINE which stands for “Autoregressive Learning for Planning In NEtworks.” Project ALPINE initiates a theoretical investigation into the development of planning capabilities in Transformer-based language models through their autoregressive learning mechanisms, aiming to identify any potential limitations in their planning abilities. We abstract planning as a network path-finding task where the objective is to generate a valid path from a specified source node to a designated target node. In terms of expressiveness, we show that the Transformer is capable of executing path-finding by embedding the adjacency and reachability matrices within its weights. Our theoretical analysis of the gradient-based learning dynamic of the Transformer reveals that the Transformer is capable of learning both the adjacency matrix and a limited form of the reachability matrix. These theoretical insights are then validated through experiments, which demonstrate that the Transformer indeed learns the adjacency matrix and an incomplete reachability matrix, which aligns with the predictions made in our theoretical analysis. Additionally, when applying our methodology to a real-world planning benchmark, called Blocksworld, our observations remain consistent. Our theoretical and empirical analyses further unveil a potential limitation of Transformer in path-finding: it cannot identify reachability relationships through transitivity, and thus would fail when path concatenation is needed to generate a path. In summary, our findings shed new light on how the internal mechanisms of autoregressive learning enable planning in networks. This study may contribute to our understanding of the general planning capabilities in other related domains.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, have impressed everyone with their powerful capabilities in multi-faceted tasks spanning language processing, knowledge extraction, reasoning, planning, coding, tool use, and more. The broad spectrum of intelligent capabilities exhibited by LLMs reflects promising signs of artificial general intelligence (AGI) [BCE+23] and catalyzes an AI revolution: Individuals and organizations are now striving to develop more powerful and adaptive AI models towards AGI, while also integrating LLM-based AI models into various aspects of our work and daily lives. However, at the same time, we are still intrigued by the underlying mechanisms that fuel the power of LLMs. While all current LLMs are built upon the Transformer neural network architecture, which employs autoregressive learning to predict the next word in a language sequence, the question remains: Why does the Transformer-based autoregressive learning architecture produce such exceptional performance across a wide range of intelligent tasks? To put it in plain English: Why does next word prediction generate intelligence?
There is no definite answer to this question yet. But researchers are tackling this problem from various angles, aiming to provide explanations to the power of LLMs. In this paper, we focus on the planning capability of LLMs. Planning is a fundamental construct of human intelligence and is involved in almost every aspect of our daily life, e.g., planning a task at work, organizing a trip, seeking a mathematical proof of a theorem, etc. Understanding how LLMs completes a planning task can shed light on the transformation of the seemingly low-level statistical task of next word prediction into a high-level intelligent task like planning. This understanding may serve as a potential pathway to comprehend and explain other intelligent behaviors exhibited by LLMs. There are already a number of studies on the planning capabilities of LLMs. But most of them mainly focus on the empirical evaluation of LLMs. Although these studies may show some evidences that LLMs have planning capabilities to a limited extent, the results are partial and do not explain why LLMs can or cannot successfully accomplish specific planning tasks (see Section 7 for more detailed discussions on the related work).
In light of this context, we initiated Project ALPINE, which stands for Autoregressive Learning for Planning In NEtworks. Project ALPINE aims to not only empirically evaluate the planning performance of LLMs but also provide theoretical interpretations on how LLMs accomplish such tasks. To provide a solid foundation for our investigation, we must define a specific task that serves as a representative example of a planning task. Given that planning often entails making sequential selections of next steps within a multi-step procedure to achieve a desired goal, it naturally relates to the path-finding task in networks.
A complex task is often represented as a task graph, where nodes correspond to subtasks or intermediate stages, and edges represent the ordering relationships betweeb these subtasks. Task planning involves finding a valid path within the task graph to reach a pre-determined goal. For example, project planning can be viewed as navigating through the multi-stage project planning graph, while a mathematical proof can be seeing as a path from the axioms to the final theorem, with lemmas serving as intermediate nodes. Many previous studies on LLM planning capabilities are also related to path finding. For example, a benchmark planning game called Blocksworld [VMSK24] used for evaluating LLMs can be viewed as path finding from the initial blocks’ state to the final blocks’ state in a state transition graph. Similarly, HuggingGPT [SST+23] for scheduling API calls can be likened to finding a call path in the API call graph.
In Project ALPINE, we investigate the following path-finding task: given an underlying graph, the training data consists of a collection of paths in the graph that specify the source node $s$, the target node $t$ and a path from $s$ to $t$. The objective of the test is to generate a path from $s^{\prime}$ to $t^{\prime}$, given new source-target pair $(s^{\prime},t^{\prime})$.
Note that, for this path-finding task, the generative model must generate a valid path in one shot without relying on trial and error. In this case, the key challenge, when giving the current node on the path, lies in correctly identifying the next node on the path, and this node should be both adjacent to the current node and reachable to the target node (see Algorithm 1). This suggests that in order to accomplish the path-finding task, it is essential to extract the information about the adjacency and reachability of the graph from the training data. Our investigation below demonstrates that the Transformer model is indeed performing this extraction to a certain extent.
More specifically, we investigate how the Transformer-based autoregressive learning architecture achieves the path-finding task by examining the following aspects. First, we show that the Transformer architecture has the expressive power to complete the task by manually constructing a Transformer that encodes the adjacency matrix and reachability matrix of the network as part of its model. Second, we conduct theoretical analysis to further explore the capabilities of the Transformer model. For a simplified Transformer model, when applying gradient descent to minimize the cross-entropy loss on the path training data, our analysis reveals that the model can extract the adjacency matrix and a limited form of the reachability matrix. Specifically, the feed-forward layer encodes the adjacency matrix, while the attention layer captures a partial form of the reachability matrix. This process mimics human intelligence in generating the next node that is both adjacent to the current node and reachable to the target node. However, our theoretical analysis also reveals a limitation of the Transformer model. It cannot learn the full complete reachability relationship. Particularly the reachability derived from transitivity cannot be learned by the model as the Transformer falls short in capturing these complex reachability patterns.
Third, we conduct extensive experiments that train Transformer models on the path language through autoregressive learning, and test its capability for generating valid paths for new pairs of source and target nodes. Our empirical results provide compelling evidence that the Transformer can excel in achieving high accuracy in the path-finding task. These findings also align with our theoretical analysis as they show that the trained Transformer model generates the next node on the path by focusing its attention on the target node, and effectively learns the adjacency and reachability matrices in its feed-forward layer and attention layer, respectively. Moreover, we observe a significant drop in test accuracy when the source and target can only be connected through concatenation of path segments in the training data. This indicates the requirement for higher-order transitive relationship to establish reachability. This matches our theoretical analysis, showing that the Transformer indeed has limitation in learning transitive reachability relationships. Finally, we demonstrate that the Transformer can successfully learn a task planning benchmark called Blocksworld [VMSK24], a planning task that corresponds directly to the path-finding problem.
In summary, our investigation in Project ALPINE makes the following contributions: (a) We have initiated a theoretical analysis that explains how Transformer achieves a path-planning task through its gradient descent autoregressive learning mechanism; (b) Our empirical evaluation corroborates our theoretical analysis and clearly demonstrates how the Transformer accomplishes path planning by extracting the adjacency and reachability information, while focusing attention on the target node. (c) Both our theoretical and empirical analyses uncover a limitation of the Transformer architecture, highlighting its inability to handle transitive reachability relationship in the path-finding task, which holds significant implications. Our findings represent an initial step toward unraveling the underlying mechanism that empowers the intelligent of LLMs. We believe that this meaningful first step brings us closer to achieving our ultimate objective. We hope that our findings, along with our integrated theoretical and empirical approach, will prove valuable to the community, facilitating collective progress in understanding LLMs and driving improvements future generations of these models.
The rest of the paper is organized as follows. In Section 2, we provide the preliminaries for the study. Section 3 presents an overview of our technical results, including the expressive power of the Transformer model in the path-finding task. In Section 4, we present our theoretical analysis on how a simplified Transformer model solves the path-finding task. Section 5 provides the detailed empirical evaluation results that reinforce our theoretical analysis. Section 6 highlights our findings on the Blocksworld task, which are consistent with our main findings. Following that, in Section 7, we provide an overview of related work and discuss the implications of our results. Finally, we conclude the paper in Section 8.
@@ 2 Preliminaries
Throughout this paper, we use the following notations for matrices and vectors: $\bm{a}$ and $\bm{A}$ stand for a column vector and a matrix, respectively. Notations $\bm{a}_{(i)}$ and $\bm{A}_{(i,j)}$ denote the $i^{th}$ entry of vector $\bm{a}$ and the $(i,j)^{th}$ entry in matrix $\bm{A}$, respectively. We also denote the $i^{th}$ row of matrix $\bm{A}$ by $\bm{A}_{(i,:)}$ and the transpose of $\bm{A}$ by $\bm{A}^{\top}$.
2.1 Auto-regressive Transformer Architecture and Loss Function
@@
In this paper, we adopt the NanoGPT architecture. Let $N$ denote the sequence length, $d$ the embedding size, $H$ the number of heads, $d_{k}=d/H$ the embedding size per head, and $M$ the vocabulary size.
One key component of the architecture is its attention mechanism, which is formulated as
@@
$\text{Attention}(\bm{Q},\bm{K},\bm{V})=\textbf{softmax}\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d_{k}}}\right)\bm{V},$ (1)
where $\bm{Q}\in\mathbb{R}^{N\times d_{k}}$, $\bm{K}\in\mathbb{R}^{N\times d_{k}}$, $\bm{V}\in\mathbb{R}^{N\times d_{k}}$ are the query, key, and value matrices, respectively. Function softmax takes a vector $\bm{a}\in\mathbb{R}^{m}$ and transforms it into $\bm{b}\in\mathbb{R}^{m}$ with $\bm{b}_{i}=\frac{e^{\bm{a}_{i}}}{\sum_{j=1}^{m}e^{\bm{a}_{j}}}$. When softmax applies to a matrix, it applies to every row of the matrix.
Denoting $\bm{X}\in\mathbb{R}^{N\times d}$ as input, the multi-head attention is computed as
$\text{MHA}(\bm{X})=\text{Concat}_{i\in[H]}(\text{Attention}(\bm{X}\bm{W}_{i}^{Q},\bm{X}\bm{W}_{i}^{K},\bm{X}\bm{W}_{i}^{V})),$ (2)
@@
where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}^{K}_{i}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}^{V}_{i}\in\mathbb{R}^{d\times d_{k}}$ are the learnable weight matrices for the query, key, and value matrices of the $i^{th}$ head.
The feed-forward layer is a two-layer multi-layer perceptron (MLP) defined as
$\text{FFN}(\bm{X})=\max(\bm{0},\bm{X}\bm{W}_{1}+\bm{1}_{N\times 1}\bm{b}_{1}^{\top})\bm{W}_{2}+\bm{1}_{N\times 1}\bm{b}_{2}^{\top},$ (3)
@@
where $\bm{W}_{1}\in\mathbb{R}^{d\times 4d}$, $\bm{W}_{2}\in\mathbb{R}^{4d\times d}$, $\bm{b}_{1}\in\mathbb{R}^{4d}$, and $\bm{b}_{2}\in\mathbb{R}^{d}$ are the learnable weight matrices of FFN, and $\bm{1}_{N\times x}$ denotes the all-one matrix with dimension $N\times x$. Finally, one-layer Transformer is defined as
$\text{Transformer}(\bm{X})=\text{FFN}(\text{LN}_{2}(\text{MHA}(\text{LN}_{1}(\bm{X}))+\bm{X}))+\text{MHA}(\text{LN}_{1}(\bm{X}))+\bm{X},$ (4)
where $\text{LN}_{1}$ and $\text{LN}_{2}$ are two layer normalizations. The layer normalization is defined as $\text{LN}_{i}(\bm{X})=\frac{\bm{X}-\mu}{\sigma+\epsilon}\cdot\gamma_{i}+\beta_{i}$, where the expectation $\mu\in\mathbb{R}$ and standard deviation $\sigma\in\mathbb{R}$ are averaged across all terms in $\bm{X}$, $\epsilon=10^{-5}$, and $\gamma_{i},\beta_{i}$ are two learnable scalars.
@@
With these essential components in place, we proceed to introduce the procedures of GPT. The training data consists of a sequence of tokens $\bm{u}=(u_{1},\cdots,u_{N})$, where $u_{n}$ is the token id for the $n^{th}$ token. We first represent the tokens by the one-hot embedding matrix $\bm{U}\in\mathbb{R}^{N\times M}$, where $\bm{U}_{(i,u_{i})}=1$ and 00 elsewhere. Then there is a learnable token embedding matrix $\bm{W}_{t}\in\mathbb{R}^{M\times d}$ and positional embedding matrix $\bm{W}_{p}\in\mathbb{R}^{N\times d}$, and the input $\bm{H}_{0}=\bm{U}\bm{W}_{t}+\bm{W}_{p}\in\mathbb{R}^{N\times d}$.
This input $\bm{H}_{0}$ is fed into an $L$-layer Transformer to obtain the predicted next word [^1] The learnable weight matrices of different layers are different, and thus the layer index $l$ should be added as a subscript to these matrices. But our later analysis is only on a one-layer Transformer. Thus we omitted this layer index.:
$\bm{H}_{l}=\text{Transformer}(\bm{H}_{l-1})\in\mathbb{R}^{N\times d},l=1,\cdots,L.$ (5)
@@
Finally, the output embedding goes through another layer normalization $\text{LN}_{t}$, and then it is multiplied by a learnable output weight matrix $\bm{W}_{o}\in\mathbb{R}^{d\times M}$ to convert back to probability weights over all possible tokens.
We calculate the output probability vector at position $n$, denoted as $\hat{\bm{u}}_{(n+1)}$, to predict the next token for position $n+1$, which reflects the auto-regressive learning paradigm:
$\hat{\bm{u}}_{(n+1)}=\text{softmax}((\text{LN}_{t}(\bm{H}_{L}))_{(n,:)}\bm{W}_{o}),1\leq n<N.$ (6)
The actual token $u_{k+1}$ is sampled according to the probability vector $\hat{\bm{u}}_{(n+1)}$ and a temperature parameter. When temperature parameter is set to $1$, which is what we use throughout the paper, the sample is directly proportionally to the probability value in $\hat{\bm{u}}_{(n+1)}$.
@@
The adopted loss function is the cross-entropy loss for the next token prediction given by
$\ell(\bm{U})=-\sum_{n=1}^{N-1}\sum_{k=1}^{M}\bm{U}_{(n+1,k)}\log\hat{\bm{u}}_{(n+1),k}.$ (7)
2.2 Path Planning Dataset

### v3

@@ Abstract
Planning is a crucial element of both human intelligence and contemporary large language models (LLMs). In this paper, we initiate a theoretical investigation into the emergence of planning capabilities in Transformer-based LLMs via their next-word prediction mechanisms. We model planning as a network path-finding task, where the objective is to generate a valid path from a specified source node to a designated target node. Our mathematical characterization shows that Transformer architectures can execute path-finding by embedding the adjacency and reachability matrices within their weights. Furthermore, our theoretical analysis of gradient-based learning dynamics reveals that LLMs can learn both the adjacency and a limited form of the reachability matrices. These theoretical insights are then validated through experiments, which demonstrate that Transformer architectures indeed learn the adjacency and an incomplete reachability matrices, consistent with our theoretical predictions. When applying our methodology to the real-world planning benchmark Blocksworld, our observations remain consistent. Additionally, our analyses uncover a fundamental limitation of current Transformer architectures in path-finding: these architectures cannot identify reachability relationships through transitivity, which leads to failures in generating paths when concatenation is required. These findings provide new insights into how the internal mechanisms of autoregressive learning facilitate intelligent planning and deepen our understanding of how future LLMs might achieve more advanced and general planning-and-reasoning capabilities across diverse applications.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT have impressed many with their powerful capabilities across a wide range of tasks, including language processing, knowledge extraction, reasoning, planning, coding, tool use, and more [1]. However, we continue to be intrigued by the underlying mechanisms that fuel the power of LLMs. While all current LLMs are built on the Transformer architecture, which uses autoregressive learning to predict the next word in a language sequence, the overarching question remains:
Why does the process of next-word prediction give rise to intelligence?
There is no definite answer to this question yet, but researchers are approaching the problem from various angles, aiming to characterize the power and limitations of LLMs, as well as to capture their underlying acquisition, abstraction, generalization, adaptation, and reasoning mechanisms. Recently, the mechanisms of grammar learning, knowledge manipulation, scaling laws, and arithmetic operations have been empirically uncovered [4, 3, 5, 2, 31, 11]. Furthermore, theoretical analyses have been conducted on in-context learning, chain-of-thought, and other forms of reasoning [30, 8, 7, 27]. Beyond these, LLMs’ capability for planning—a fundamental component of human intelligence—has also drawn considerable attention. Planning is involved in nearly every aspect of our daily life, such as organizing a task at work, planning a trip, or seeking a mathematical proof of a theorem. Additionally, task planning plays a pivotal role in state-of-the-art LLM-empowered autonomous agents, such as HuggingGPT [18], Voyager [23], and Reflection [19]. Understanding how LLMs complete a planning task can shed light on how the seemingly low-level statistical task of next-word prediction transforms into a high-level intelligent process. Several prior studies have empirically evaluated the planning capabilities of LLMs, yielding both positive and negative results [16, 22]. However, the current results are incomplete and do not fully explain why LLMs can or cannot successfully accomplish specific planning tasks (see Appendix A for a detailed discussion of related works).
Given that planning often involves making sequential selections of next steps to achieve a desired goal, it naturally relates to the path-finding task in networks. For example, autonomous agents (e.g., HuggingGPT [18]) for scheduling API calls can be likened to finding a call path in the API call graph; a mathematical proof can be seen as a proof path from the axioms to the final theorem [21]; and a step-by-step solution to a grade-school math problem can be viewed as a path in the dependency graph among the variables [28, 29]. Many previous studies on LLM planning capabilities are related to path-finding. e.g., an LLM planning benchmark called Blocksworld [22] can be viewed as path-finding from the initial state of the blocks to the final state in a state transition graph. Furthermore, in neuroscience, planning is often evaluated through path-finding in a maze [26]. Consequently, in this paper, we abstract planning in LLM learning as the following path-finding task: given an underlying directed graph, a Transformer architecture is provided with training data consisting of a collection of paths that specify the source node, the target node, and a path from the source to the target. The task of the language model is then to generate a path for a new source-target pair. In addition to measuring the performance of the trained model, we examine the internal weighting mechanism and the learning dynamics of the Transformer architecture during the learning and problem-solving process. This research is part of our broader project, ALPINE (Autoregressive Learning for Planning In NEtworks), which aims to answer the overarching question on the connection between the process of next-word prediction and the emergence of intelligence through the lens of planning.
Our Contributions: Our project initiates a theoretical investigation into the development of planning capabilities in Transformer-based language models by focusing on characterizing their expressiveness and learning dynamics in the path-finding task. First, in Theorem 2, we present a mathematical construction of a Transformer that encodes both the adjacency and reachability matrices of the network, thereby establishing that Transformer architectures possess the expressive capacity to complete the path-finding task. Then, in Theorem 3, we prove that when applying gradient descent to minimize the cross-entropy loss on the training data, a model based on a simplified Transformer architecture can extract the adjacency matrix and a limited form of the reachability matrix, using them to mimic human-like intelligence in path-finding. Our theoretical analysis further reveals a fundamental limitation of current Transformer architectures: they do not learn certain types of reachability, particularly transitive reachability, resulting in an incomplete ability to reason about future steps when planning. To validate our theoretical findings, we conduct extensive experiments training Transformers on the path language using autoregressive learning. First, these experiments demonstrate that Transformers achieve high accuracy in the path-finding task (Figure 3). Second, we show that it is indeed possible to extract both the adjacency and a limited form of the reachability matrices from the Transformers’ weights (Figures 1,2,5,6(a)). Third, we observe a significant drop in test accuracy when the source and target nodes are connected only through concatenated path segments in the training data (Figure 6). These findings align with our theoretical analysis, confirming that current Transformers have limitations in learning transitive reachability relationships, unlike human intelligence. Finally, we validate these results on a real-world task planning benchmark, Blocksworld [22], which directly corresponds to the path-finding problem (see Appendix F).
@@ 2 Preliminaries
Throughout this paper, we use the following notations for matrices and vectors: $\bm{a}$ and $\bm{A}$ stand for a column vector and a matrix, respectively. Notations $\bm{a}_{(i)}$ and $\bm{A}_{(i,j)}$ denote the $i^{th}$ entry of vector $\bm{a}$ and the $(i,j)^{th}$ entry in matrix $\bm{A}$, respectively. We also denote the $i^{th}$ row of matrix $\bm{A}$ by $\bm{A}_{(i,:)}$.
2.1 Autoregressive Transformer Architecture and Loss Function
@@
In this paper, we adopt the standard GPT architecture [17]. We use the following notation for the architecture and loss function in our analysis. Let $N$ denote the sequence length, $d$ the embedding size, $H$ the number of heads, $d_{k}=d/H$ the embedding size per head, and $M$ the vocabulary size. One key component of the architecture is the attention mechanism, which is formulated as:
@@
$\text{Attention}(\bm{Q},\bm{K},\bm{V})=\textbf{softmax}\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d_{k}}}\right)\bm{V}$ (1)
@@
where $\bm{Q}\in\mathbb{R}^{N\times d_{k}}$, $\bm{K}\in\mathbb{R}^{N\times d_{k}}$, $\bm{V}\in\mathbb{R}^{N\times d_{k}}$ are the query, key, and value matrices, respectively. Denoting $\bm{X}\in\mathbb{R}^{N\times d}$ as input, the multi-head attention is computed as $\text{MHA}(\bm{X})=\text{Concat}_{i\in[H]}(\text{Attention}(\bm{X}\bm{W}_{i}^{Q},\bm{X}\bm{W}_{i}^{K},\bm{X}\bm{W}_{i}^{V}))$, where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}^{K}_{i}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}^{V}_{i}\in\mathbb{R}^{d\times d_{k}}$ are the learnable weight matrices for the query, key, and value matrices of the $i$-th head.
The feed-forward layer is a two-layer multi-layer perceptron (MLP) defined as follows:
$\text{FFN}(\bm{X})=\max(\bm{0},\bm{X}\bm{W}_{1}+\bm{1}_{N\times 1}\bm{b}_{1}^{\top})\bm{W}_{2}+\bm{1}_{N\times 1}\bm{b}_{2}^{\top}$ (2)
@@
where $\bm{W}_{1}\in\mathbb{R}^{d\times 4d}$, $\bm{W}_{2}\in\mathbb{R}^{4d\times d}$, $\bm{b}_{1}\in\mathbb{R}^{4d}$, and $\bm{b}_{2}\in\mathbb{R}^{d}$ are the learnable weight matrices and $\bm{1}_{N\times x}$ is the all-one matrix with dimension $N\times x$. Finally, one Transformer layer is defined as:
$\text{Transformer}(\bm{X})=\text{FFN}(\text{LN}_{2}(\text{MHA}(\text{LN}_{1}(\bm{X}))+\bm{X}))+\text{MHA}(\text{LN}_{1}(\bm{X}))+\bm{X}$ (3)
where $\text{LN}_{1}$ and $\text{LN}_{2}$ are two layer normalizations.
@@
With these essential components in place, we proceed to introduce the procedures of GPT. The training data consists of many sequences of tokens, where each sequence is expressed as $\bm{u}=(u_{1},\cdots,u_{N})$, in which $u_{n}$ denotes the token id for the $n$-th token in sequence $\bm{u}$. We first represent the tokens in $\bm{u}$ by a one-hot embedding matrix $\bm{U}\in\mathbb{R}^{N\times M}$, where $\bm{U}_{(n,u_{n})}=1$ and 00 elsewhere. Then there are learnable token embedding matrix $\bm{W}_{t}\in\mathbb{R}^{M\times d}$ and positional embedding matrix $\bm{W}_{p}\in\mathbb{R}^{N\times d}$, and the input $\bm{H}_{0}=\bm{U}\bm{W}_{t}+\bm{W}_{p}\in\mathbb{R}^{N\times d}$. This input $\bm{H}_{0}$ is fed into an $L$-layer Transformer, i.e., $\bm{H}_{l}=\text{Transformer}(\bm{H}_{l-1})\in\mathbb{R}^{N\times d}$ for $l=1,\cdots,L$.
@@
Finally, the output embedding goes through another layer normalization $\text{LN}_{t}$, and then it is multiplied by a learnable output weight matrix $\bm{W}_{o}\in\mathbb{R}^{d\times M}$ to convert back to probability weights over all possible tokens. We calculate the output probability vector at position $n$, denoted by $\hat{\bm{u}}_{(n+1)}$, using $\hat{\bm{u}}_{(n+1)}=\textbf{softmax}((\text{LN}_{t}(\bm{H}_{L}))_{(n,:)}\bm{W}_{o}),1\leq n<N$. This probability vector is used to predict the token for position $n+1$, which reflects the autoregressive learning paradigm.
@@
The adopted loss function is the cross-entropy loss for the next token prediction, given by:
$\ell(\bm{u})=-\sum_{n=1}^{N-1}\sum_{k=1}^{M}\bm{U}_{(n+1,k)}\log\hat{\bm{u}}_{(n+1),k}$ (4)
2.2 Path-Planning Dataset: Syntax and Data Sources